OpenStudy (anonymous):

I need help with factoring a quadratic equation with only two terms. The equation is 0=-16t^2+230. I have to factor it but my course never taught me how to factor with only two terms. Please help!

OpenStudy (anonymous):

ah try this? 16t^2 + Ot -230 = 0 now solve

OpenStudy (anonymous):

How do I factor it with the zero? I'm supposed to split it into two binomial terms.

OpenStudy (anonymous):

ofcrs you can think

OpenStudy (anonymous):

simple method 16t^2 = 230 t^2 = 230/16 solve n you will get 2 values

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

I have t= 7.187

OpenStudy (anonymous):

Is the second one t= -7.187?

OpenStudy (anonymous):

idk,i didnt solve but you are supposed to factorize,so ur teachers will tell u my first method,wait for them to explain